OpenStudy (anonymous):

A tough one. Find the largest possible area for A(theta)=49sin(2theta)

OpenStudy (anonymous):

\[A(\Theta)=49\sin 2(\Theta)\]

OpenStudy (amistre64):

you know how to do derivatives?

OpenStudy (anonymous):

haha some but not with trigonometric functions. I am in precalc

OpenStudy (amistre64):

ok...lets try this route; the largest area of any "sin" is a full circle right?

OpenStudy (anonymous):

yes, 2pi

OpenStudy (amistre64):

then we need to find a circle that has the same "information" is the given equation.

OpenStudy (amistre64):

Do you know the area of a sector? when it involves 2 sides and the sin of an angle?

OpenStudy (anonymous):

not yet...I think thats in an upcoming chapter.

OpenStudy (amistre64):

ok... and maybe I saw this wrong to begin with...consider this: the area of a triangle is: (1/2) (base) (height) right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

How do we find the height of a triangle when we know an angle? does sin(t) = y/r?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

like this

OpenStudy (anonymous):

yes...I understand that

OpenStudy (amistre64):

good, then we need to determine the value of "t" that will give us the largest area for the triangle....or thats how I understand this problem

OpenStudy (anonymous):

and the biggest sin you can get is 1

OpenStudy (anonymous):

would it help to tell you that I have a rectangle inscribed inside a semicircle with a radius of 7cm.

OpenStudy (anonymous):

and that is the equation for the area

OpenStudy (amistre64):

i keep seeing this 2 different ways...for example: the area of any given sector of a circle is define as (theta)(r^2)/2

OpenStudy (amistre64):

it would..... all the information helps :)

OpenStudy (anonymous):

sorry about that

OpenStudy (amistre64):

like this?

OpenStudy (amistre64):

like this?

OpenStudy (anonymous):

exactly!

OpenStudy (amistre64):

yeah....thatd a been helpful to know lol

OpenStudy (anonymous):

haha...wow...sorry

OpenStudy (amistre64):

which angle is the theta...near the center of up near a corner?

OpenStudy (anonymous):

near the center

OpenStudy (amistre64):

if I recall correctly, the angle needs to be 45 degrees. so 2t = 45 t = 22.5

OpenStudy (amistre64):

the largest area you can get is formed by a square.... and you need 2 squares side by side to make this the max area under the half circle

OpenStudy (anonymous):

so would my dimension them be 7 by 14?

OpenStudy (amistre64):

2 sides equal 7sin(45); and the other 2 sides are 28sin(45) 4.95 is one side... 9.90 is the other side...... it gives you an area about equal to 49.005

OpenStudy (anonymous):

how did you get these?

OpenStudy (amistre64):

magic :) first I killed a live chicken; then a spread its blood around the bedposts......

OpenStudy (anonymous):

haha seriosuly!...thats how I feel about these problems! You need to do voodoo to get an answer!!

OpenStudy (amistre64):

we take the equation for a half circle: y = sqrt(7^2-x^2) and the Area of the "rectangle" to be maximized. A = xy ; substitute y = sqrt(49 - x^2) into this equation

OpenStudy (amistre64):

A = x(sqrt(49-x^2)) now find the derivative.... A' = x(-x/(sqrt(49-x^2))) +1(sqrt(49-x^2)) and make that equal to zero.... then solve for x :)

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

its calculus....but it works :)

OpenStudy (amistre64):

the other way is just to notice that the largest area that can be produced in a quarter of a circle is a square....a square has a 45 degree angle where we need it down there

OpenStudy (anonymous):

gotcha!!

OpenStudy (anonymous):

thanks! have a good night!

OpenStudy (amistre64):

youre welcome :) Gnite